OpenStudy (anonymous):

Find the product: (6t + 4)(-5t – 10)

OpenStudy (anonymous):

refer http://adf.ly/QNJWy

OpenStudy (anonymous):

\[(6t+4)(-5t-10)=6t\times5t+6t\times(-10)+4\times(-5t)+4\times(-10) \\=30t^2-6t-20t-40 \\=30t^2-26t-40 \]